OpenStudy (anonymous):

A purse contains $4.70 in nickels and quarters. There are 30 coins in total. how many nickels and quarters are there?

OpenStudy (texaschic101):

n + q = 30 .05n + .25q = 4.70 now it can be solved by substitution. Can you do it now ?

OpenStudy (anonymous):

Yes, thanks for the help.

OpenStudy (texaschic101):

do you have any questions about the equations ?

OpenStudy (anonymous):

Yes, since there are 2 variables what do I do to actually get the answer?

OpenStudy (texaschic101):

substitution.. n + q = 30....n = 30 - q so sub in 30 - q for n back into the second equation. This will allow you to solve for q. You can then sub that answer back into the first equation and solve for n

OpenStudy (anonymous):

oh, okay. Thanks.

OpenStudy (texaschic101):

sure thing :)